8) What is investment in a closed economy if you have the following economic data? Y = $10 trillion C = $5 trillion TR = $2 trillion G = $2 trillion

a. $5 trillion b. $2 trillion c. $3 trillion d. cannot be determined without information on taxes (T)
Answer #1

GDP is the sum of consumption investment and government expenditure in the closed economy

Y = C + I + G

I = Y - C - G

I = 10 - 5 - 2 = 3 trillion

Thus, investment is 3 trillion dollar.

Select option C.